OpenStudy (anonymous):

Find two digit number( or the sum of those numbers) in which when changing the numbers, the number will be decreased by 28,125%.

OpenStudy (experimentx):

((10x+y) - (10y+x))/(10x+y) = 0.28125

OpenStudy (experimentx):

looks like i should go for programming solution than algebraic solution

OpenStudy (anonymous):

Algebraic would be needed.

OpenStudy (asnaseer):

decrease by 28.125% means new number will be 575/800 times old number. so, using experimentX's symbols:\[\frac{575}{800}(10x+y)=10y+x\]multiplying out and simplifying will lead to:\[y=\frac{2}{3}x\]you should be able to now work out all the required values.

OpenStudy (anonymous):

Thanks!

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

vw